3 5/6 - 5 2/3 and reduce to lowest
Answer #1

3 5/6 =( 6*3 + 5 )/6 = 23/6

5 2/3 =( 3*5+ 2)/3 = 17/3

23/6-17/3 = (23-34)/6

= - 11/6 = - 1 5/6